# LLMBind: A Unified Modality-Task Integration Framework

### v5

@@ 3.4. Dataset
The training datasets consist of both publicly available datasets and our custom proprietary ones, as visualized in Fig. 5. These proprietary datasets are specifically constructed through interactive processes enabled by ChatGPT.
@@
Interactive Image Generation and Editing Dataset. To empower the model in executing image generation and editing tasks, following Mini-DALLE3 (Zeqiang et al., 2023), we design a prompt for ChatGPT to construct 400k human-like interactive instruction data, tackling the problem of image generation and editing through textual description. As shown in Fig. 6, we first clarify the role and capability of ChatGPT within the prompt. We provide it with instruction text formats, tags, and examples for both generating and editing tasks. However, we notice that the generated samples closely resemble the examples in terms of text format, tone, and content. To address this, we further clarify the diversity of the target data in terms of the above multiple aspects in the prompt. Following a manual review of a portion of our dataset, we have confirmed that this approach is highly effective in ensuring significant diversity.
@@
Interactive Video and Audio Generation Dataset. To construct video and audio generation instruction data, we adapt the t2x(Wu et al., 2023) dataset and change its format to resemble that used for the image generation task. Some examples are listed in Appendix C.
@@
Segmentation Datasets. For segmentation task, we incorporate several semantic segmentation datasets, including ADE20K (Zhou et al., 2017), COCO-Stuff (Caesar et al., 2018), PACO-LVIS (Ramanathan et al., 2023), PartImageNet (He et al., 2022), and PASCAL-Part (Chen et al., 2014). We also utilize referring segmentation datasets like refCLEF, refCOCO, refCOCO+(Kazemzadeh et al., 2014), and refCOCOg(Mao et al., 2016).
Visual Question Answering Datasets. We incorporates the LLaVA-Instruct-665k dataset (Liu et al., 2023d) for visual question-answering task. This dataset enables the training of models to comprehend, reason, and make decisions accurately based on visual content.
4. Experiments
4.1. Experimental Setting
@@
Evaluation Metrics. Our experiments utilize some key metrics for precise evaluation. Text-to-audio generation is assessed using Fréchet Distance (FD) for distribution similarity and Inception Score (IS) for quality and diversity. Text-to-video generation leverages Fréchet Inception Distance (FID) for content quality and CLIPSIM for textual alignment. In text-to-image generation, FID evaluates image distribution similarity. For segmentation, gIoU is defined by the average of all per-image Intersection-over-Unions (IoUs), while cIoU is defined by the cumulative intersection over the cumulative union. In addition, to better showcase the model’s generative capabilities during user interaction, we employed human evaluation to assess the quality of the modality data generated and its alignment with the given text instructions.
4.2. Results Visualization
As shown in Fig. 7, LLMBind can be executed on various modal tasks such as video, audio, image generation, image editing, and image segmentation, providing users with a more user-friendly interactive experience.
4.3. Quantitative Analysis
@@
Referring Segmentation. For referring segmentation task, as illustrated in Table 2, LLMBind achieves superior results on various datasets. On the refCOCO dataset, for the val, testA, and testB splits, LLMBind achieves scores of 76.9, 78.5, and 73.2, respectively. Similarly, on the refCOCO dataset, the model records scores of 67.8, 71.9, and 60.8 for val, testA, and testB. On the refCOCOG dataset, LLMBind achieves 69.8 and 70.8.
@@
Image Generation. In the text-to-image generation task, as shown in Table 3, LLMBind utilizes various baseline models, achieving impressive FID scores on the COCO-caption (Lin et al., 2014) dataset. Specifically, with Stable Diffusion v1.5 and v2.0 (Rombach et al., 2022c) as baselines, LLMBind attains FID scores of 16.41 and 14.29, respectively. Notably, using GALIP (Tao et al., 2023) as the baseline model, LLMBind secures an FID score of 10.38, outperforming models such as NeXT-GPT (Wu et al., 2023) and CoDi (Tang et al., 2023).
Audio Generation. In the domain of text-to-audio generation, LLMBind has been meticulously evaluated and compared with existing models. As depicted in Table 5, LLMBind utilizes the CoDi (Tang et al., 2023) model as a baseline for audio generation, attaining an FD score of 22.90 and an IS of 8.77 on the AudioCaps (Kim et al., 2019) dataset. Furthermore, when employing Auffusion (Tang et al., 2023) as the foundational model for generation, LLMBind secures an FD score of 21.99 and an IS score of 10.57 on the same dataset, demonstrating a discernible advantage over models such as NeXT-GPT (Wu et al., 2023) and AudioLDM (Liu et al., 2023a).
| Method | Q.S. | Final | A.S. | Final | ||||||
| 1 | 2 | 3 | 4 | score | 1 | 2 | 3 | 4 | Score | |
| Audio Generation | ||||||||||
| Direct input | 7% | 35% | 34% | 24% | 2.75 | 18% | 39% | 23% | 20% | 2.45 |
| Ours | 8% | 28% | 35% | 29% | 2.85 | 6% | 24% | 27% | 43% | 3.07 |
| Image Generation | ||||||||||
| Direct input | 5% | 39% | 43% | 13% | 2.64 | 21% | 29% | 45% | 5% | 2.34 |
| Ours | 5% | 27% | 41% | 27% | 2.90 | 12% | 19% | 51% | 18% | 2.75 |
| Video Generation | ||||||||||
| Direct input | 29% | 41% | 23% | 7% | 2.08 | 23% | 42% | 21% | 14% | 2.26 |
| Ours | 24% | 38% | 27% | 11% | 2.25 | 19% | 40% | 24% | 17% | 2.39 |
Video Generation. For the text-to-video generation task, LLMBind’s performance on the MSR-VTT (Xu et al., 2016) dataset is shown in Table 4. Using ModelScopeT2V (Wang et al., 2023) as the base model, LLMBind achieves an FID of 11.09 and a CLIPSIM score of 0.2930. With Show-1 (Zhang et al., 2023) as the base model, it obtains an FID of 13.08 and a CLIPSIM score of 0.3072. When employing HiGen (Qing et al., 2023) as the base model, LLMBind reaches an FID of 8.60 and a CLIPSIM score of 0.2947, showing comparable performance to NExT-GPT (Wu et al., 2023). As LLMBind avoids lossy training on generation models, its generative power primarily relies on the selected base models.
4.4. Human Evaluation
As shown in Fig. 8, to assess the practical performance of LLMBind in interactive Image, Video, and Audio generation tasks (IVA), we conducted a comprehensive user study. This study is designed to measure the capability of models to generate IVA data in response to interactive text instructions, emulating natural human-like communication. The study focuses on assessing two critical metrics: Quality Score (Q.S.) and Alignment Score (A.S.). The Q.S. assesses the quality of the generated IVA content, while A.S. evaluates the alignment between the generated data and the textual instructions provided. Participants interact with the model by issuing text instructions and then rate each generated sample on a scale of 1 to 4 for both metrics, with 1 representing poor performance and 4 indicating excellent performance. The aggregated results from a panel of 30 participants, as illustrated in Table 6, clearly indicate that LLMBind excels in both the quality of modal data generation and its ability to align with text-based instructions closely. The final scores, calculated as the weighted average of the ratings for Q.S. and A.S., demonstrate LLMBind’s proficiency in understanding and executing complex, multimodal tasks as directed by human users.
4.5. Ablation Analysis
We apply the ablation analysis on the reasoning segmentation task, since it is conducive to quantitative analysis and comparison. The detailed tables are listed in the supplementary.
Rank of LoRA. As shown in Table 7, our framework exhibits slight performance variation across different settings.
Training Steps. As shown in Table 8, we reveal a clear trend that longer iteration steps consistently yield improvements in both gIoU and cIoU metrics.
Number of MoE Layers. As shown in Table 9, model performance generally improves with the increase in the number of MoE layers.
Number of All Experts and Top-K Activated Experts. As shown in Table 10, when the number of experts is fixed, more activated experts improve model performance. Conversely, increasing the number of experts enhances performance when the number of activated experts is constant.
@@ 4.5. Ablation Analysis
| LoRA | Training | MoE | Expert | Top-K | gIoU | cIoU |
| RANK | Steps | Layers | Nums | Experts | ||
| 8 | 7k | 16/32 | 4 | 2 | 59.4 | 66.5 |
| 8 | 7k | 16/32 | 3 | 2 | 58.9 | 65.8 |
| 8 | 7k | 16/32 | 2 | 2 | 58.7 | 66.2 |
| 8 | 7k | 16/32 | 2 | 1 | 58.6 | 65.2 |
5. Conclusion and Future Direction
In this work, we introduce the LLMBind framework, designed to address the limitations of current Multimodal Large Language Models in handling a wide range of multi-modality tasks. LLMBind efficiently integrates tasks across various modalities such as image segmentation, video generation, and more, through its unique Task-Specific tokens. To enhance the adaptability and efficiency of our framework, we introduce a novel dataset of 400k interactive instructions and an innovative Mixture-of-Experts architecture. Extensive experimental validation demonstrates that LLMBind excels in a variety of tasks and holds immense potential for expansion to other modality tasks.
Our framework marks a significant step towards creating a unified AI agent capable of handling universal modalities. Looking ahead, we aim to further expand the LLMBind framework to include more modality tasks while concurrently focusing on investigating and refining training strategies for multi-task joint training.
@@ Appendix A Interactive Architecture
A.1. Multi-Turn and Interactive
@@
A.2. Interactive Multi-modal Generation
@@
A.3. Interactive Image Editing and Refinement
@@
A.4. Interactive Text Conversation

### v6

@@ 3.4 Dataset
The training datasets consist of both publicly available datasets and our custom proprietary ones, as visualized in Figure 4. These proprietary datasets are specifically constructed through interactive processes enabled by ChatGPT.
@@
Interactive Image Generation and Editing Dataset. To empower the model in executing image generation and editing tasks, following Mini-DALLE3 (Zeqiang et al., 2023), we design a prompt for ChatGPT to construct 400k human-like interactive instruction data, tackling the problem of image generation and editing through textual description. As shown in Figure 5, we first clarify the role and capability of ChatGPT within the prompt. We provide it with instruction text formats, tags, and examples for both generating and editing tasks. However, we notice that the generated samples closely resemble the examples in terms of text format, tone, and content. To address this, we further clarify the diversity of the target data in terms of the above multiple aspects in the prompt. Following a manual review of a portion of our dataset, we have confirmed that this approach is highly effective in ensuring significant diversity.
@@
Interactive Video and Audio Generation Dataset. To construct video and audio generation instruction data, we adapt the t2x(Wu et al., 2023b) dataset and change its format to resemble that used for the image generation task. Some examples are listed in Appendix C.
@@
Segmentation Datasets. We incorporate several semantic segmentation datasets including ADE20K (Zhou et al., 2017), COCO-Stuff (Caesar et al., 2018), PACO-LVIS (Ramanathan et al., 2023), PartImageNet (He et al., 2022), and PASCAL-Part (Chen et al., 2014) and utilize referring segmentation datasets like refCOCO, refCOCO+(Kazemzadeh et al., 2014), and refCOCOg(Mao et al., 2016).
4 Experiments
Method refCOCO refCOCO+ refCOCOg val testA testB val testA testB val(U) test(U) MCN (Luo et al., 2020) 62.4 64.2 59.7 50.6 55.0 44.7 49.2 49.4 VLT (Ding et al., 2021a) 67.5 70.5 65.2 56.3 61.0 50.1 55.0 57.7 CRIS (Wang et al., 2022) 70.5 73.2 66.1 62.3 68.1 53.7 59.9 60.4 LAVT (Yang et al., 2022) 72.7 75.8 68.8 62.1 68.4 55.1 61.2 62.1 ReLA (Liu et al., 2023a) 73.8 76.5 70.2 66.0 71.0 57.7 65.0 66.0 X-Decoder (Zou et al., 2023a) - - - - - - 64.6 - SEEM (Zou et al., 2023b) - - - - - - 65.7 - LISA (Lai et al., 2023) 74.9 79.1 72.3 65.1 70.8 58.1 67.9 70.6 UnifiedMLLM (Li et al., 2025) 76.3 78.8 72.7 66.4 72.4 59.1 68.0 69.6 LLMBind (ours) 76.9 78.5 73.2 67.8 71.9 60.8 69.8 70.8
4.1 Experimental Setting
@@
Evaluation Metrics. Text-to-audio generation is assessed using Fréchet Distance (FD) for distribution similarity and Inception Score (IS) for quality and diversity. Text-to-video generation leverages Fréchet Inception Distance (FID) for content quality and CLIPSIM for textual alignment. In text-to-image generation, FID evaluates image distribution similarity. For segmentation, gIoU is defined by the average of all per-image Intersection-over-Unions (IoUs), while cIoU is defined by the cumulative intersection over the cumulative union. In addition, to better showcase the model’s generative capabilities during user interaction, we conducted a human evaluation to assess the quality of the model output and its alignment with the given text instructions.
4.2 Quantitative Analysis
@@
Referring Segmentation. As illustrated in Table 2, on the refCOCO dataset, for the val, testA, and testB splits, LLMBind achieves scores of 76.9, 78.5, and 73.2, respectively. Similarly, on the refCOCO dataset, the model records scores of 67.8, 71.9, and 60.8 for val, testA, and testB. On the refCOCOG dataset, LLMBind achieves 69.8 and 70.8.
@@
Image Generation. As shown in Table 3, using GALIP (Tao et al., 2023) as the baseline model, LLMBind secures an FID score of 10.38, outperforming models such as NeXT-GPT (Wu et al., 2023b) and CoDi (Tang et al., 2023).
Method FD () IS () Specialized Models DiffSound (Yang et al., 2023b) 47.68 4.01 AudioLDM-S (Liu et al., 2023b) 29.48 6.90 AudioLDM-L (Liu et al., 2023b) 23.31 8.13 Multitask Models NExT-GPT (Wu et al., 2023b) 23.58 8.35 UnifiedMLLM (Li et al., 2025) 22.42 9.95 LLMBind-CoDi 22.90 8.77 LLMBind-Auffusion 21.99 10.57
Audio Generation. We evaluate LLMBind on AudioCaps (Kim et al., 2019) using different generative backbones. As shown in Table 4, with CoDi (Tang et al., 2023) as the baseline, LLMBind achieves an FD of 22.90 and an IS of 8.77. By switching to Auffusion (Tang et al., 2023), performance further improves to 21.99 (FD) and 10.57 (IS), consistently outperforming specialized models like AudioLDM (Liu et al., 2023b) and unified frameworks such as NExT-GPT (Wu et al., 2023b).
Method FID () CLIPSIM () Specialized Models CogVideo (Hong et al., 2022a) 23.59 0.2631 Latent-VDM (Rombach et al., 2022b) 14.25 0.2756 Latent-Shift (An et al., 2023) 15.23 0.2773 Make-A-Video (Singer et al., 2022a) 13.17 0.3049 CoDi (Tang et al., 2023) - 0.2890 Multitask Models NExT-GPT (Wu et al., 2023b) 13.04 0.3085 UnifiedMLLM (Li et al., 2025) 11.15 0.312 LLMBind-ModelScopeT2V 11.09 0.2930 LLMBind-HiGen 8.60 0.2947
Video Generation. We report LLMBind’s performance on MSR-VTT (Xu et al., 2016) across various generative backbones in Table 5. With ModelScopeT2V (Wang et al., 2023), LLMBind achieves an FID of 11.09 and CLIPSIM of 0.2930. Switching to HiGen (Qing et al., 2023), the performance further improves to a superior FID of 8.60, outperforming NExT-GPT (Wu et al., 2023b).
4.3 Human Evaluation
@@
As shown in Figure 6, to assess the practical performance of LLMBind in interactive Image, Video, and Audio generation tasks (IVA), we conducted a comprehensive user study. This study is designed to measure the capability of models to generate IVA data in response to interactive text instructions, emulating natural human-like communication. The study focuses on assessing two critical metrics: Quality Score (Q.S.) and Alignment Score (A.S.). The Q.S. assesses the quality of the generated IVA content, while A.S. evaluates the alignment between the generated data and the textual instructions provided. Participants interact with the model by issuing text instructions and then rate each generated sample on a scale of 1 to 4 for both metrics, with 1 representing poor performance and 4 indicating excellent performance. The aggregated results from a panel of 30 participants, as illustrated in Table 9, clearly indicate that LLMBind excels in both the quality of modal data generation and its ability to align with text-based instructions closely. The final scores, calculated as the weighted average of the ratings for Q.S. and A.S., demonstrate LLMBind’s proficiency in understanding and executing complex, multimodal generation tasks as directed by human users.
Method Q.S. Final A.S. Final 1 2 3 4 score 1 2 3 4 Score Audio Generation Direct input 7% 35% 34% 24% 2.75 18% 39% 23% 20% 2.45 Ours 8% 28% 35% 29% 2.85 6% 24% 27% 43% 3.07 Image Generation Direct input 5% 39% 43% 13% 2.64 21% 29% 45% 5% 2.34 Ours 5% 27% 41% 27% 2.90 12% 19% 51% 18% 2.75 Video Generation Direct input 29% 41% 23% 7% 2.08 23% 42% 21% 14% 2.26 Ours 24% 38% 27% 11% 2.25 19% 40% 24% 17% 2.39
4.4 Ablation Studies
We conduct the ablation study specifically on the reasoning segmentation task, as it provides a robust framework for quantitative analysis and allows for direct comparison with existing benchmarks. Specifically, we analyze how the distribution of MoE layers and the configuration of experts influence the model’s performance in handling heterogeneous multimodal tasks.
Effect of LoRA Rank.
We analyze the impact of the LoRA rank on expert capacity. As shown in Table 8, increasing the rank yields consistent but diminishing performance gains, indicating that a moderate rank is sufficient for effective expert specialization while maintaining parameter efficiency.
Analysis of MoE Layer Distribution. As illustrated in Table 8, we investigate the impact of the density of MoE layers within the 32-layer LLM backbone. We observe a consistent performance gain as the number of MoE layers increases. Specifically, increasing the MoE layers from 6 to 16 results in a significant improvement in both gIoU (from 58.3 to 59.9) and cIoU (from 61.8 to 65.8). This trend suggests that a higher density of MoE layers provides the network with more opportunities to decouple gradients from diverse tasks. By routing features through specialized experts at multiple stages of the transformer hierarchy, LLMBind effectively prevents the ”modality interference” mentioned in our introduction, ensuring that localization-sensitive features for segmentation are not overshadowed by the semantic gradients of generative tasks.
Effect of Expert Numbers and Routing Strategy. Table 8 explores the number of experts and the Top-K routing mechanism. When the expert count is increased from 2 to 4, we observe a steady increase in performance, with gIoU rising from 58.7 to 59.4. This confirms that a larger pool of experts provides greater capacity for modality disentanglement, allowing the model to allocate dedicated parameters to the specific nuances of pixel-level perception versus waveform-level synthesis. We compare the performance of Top-1 and Top-2 routing strategies. The results indicate that Top-2 routing (gIoU 58.7) outperforms Top-1 routing (gIoU 58.6). This suggests that for complex multimodal instructions, activating multiple experts allows for better collaborative reasoning. For instance, a task requiring both ”referring” (localization) and ”editing” (generation) benefits from the simultaneous activation of experts specialized in different feature domains. Consistent with our core philosophy, these results demonstrate that the MoE-LoRA architecture is essential for transforming a standard LLM into a robust multimodal system, maintaining high precision across diverse tasks without performance degradation.
5 Conclusion and Future Direction
In this work, we present LLMBind, a unified framework that redefines the MLLM as a versatile multimodal operating system. By strategically balancing In-Situ perception through Semantic Embedding Tokens and Ex-Situ generation via Task-Prompt Tokens, LLMBind overcomes the traditional trade-off between localization precision and integration flexibility. To resolve the inherent challenges of task interference in multi-task learning, we introduce a Task-Specific MoE-LoRA architecture, which ensures effective modality disentanglement and preserves specialized knowledge across heterogeneous domains.
Furthermore, our curated 400k Interactive Multi-turn Dataset provides a foundation for complex, human-like AI collaboration. Extensive experiments, particularly in reasoning segmentation and multimodal synthesis, demonstrate that LLMBind not only outperforms specialized models but also maintains high updateability for emerging modalities. Moving forward, we aim to extend this framework to a broader spectrum of sensory inputs and continue refining joint training strategies to further harmonize the synergy between perception and generation in generalist agents.
@@ Appendix A Interactive Architecture
A.1 Multi-Turn and Interactive
@@
A.2 Interactive Multi-modal Generation
@@
A.3 Interactive Image Editing and Refinement
@@
A.4 Interactive Text Conversation